OpenStudy (anonymous):

f(x) = 5, x< or=2 ax + b, 2 < x < 10 21, x> or = 10 is continuous for all x Find values a and b @rational

OpenStudy (rational):

For f(x) to be continuous at x=a we must have : LHL = RHL = f(a)

OpenStudy (rational):

``` 5, x <=2 ax + b, 2 < x < 10 ``` using LHL = RHL, get an equation for f(x) to be continuous at x=2

OpenStudy (anonymous):

So we can make two equations 5 = 2a +b and 21 = 10a + b?

OpenStudy (rational):

thats it!

OpenStudy (anonymous):

wonderful xD So that means I just solve by subtracting 10 a + b = 21 and 2a + b = 5 -> 8a = 16 or a = 16/8 a = 2 5 = 2(2) + b 5 = 4 + b b = 1 Am I right? this question is for 4 marks, on my exam day(18th this month)

OpenStudy (rational):

once you setup the equations, you're free to solve it using any method you like

OpenStudy (rational):

a=2, b=1 looks good!

OpenStudy (anonymous):

Awesome :) I might have 2 more question types under this section, is it too much trouble to ask?

OpenStudy (rational):

feel free to tag me in all your questions.. these are fun :)